OpenStudy (anonymous):

lim (4-|x|)/(sqrt(x)-2) x->4

ganeshie8 (ganeshie8):

Hint : when \(x\) is positive \[\large 4-x = 2^2 - \sqrt{x}^2 = (2+\sqrt{x})(2-\sqrt{x})\]

OpenStudy (anonymous):

thank you!